the everton college store paid $1683 for an order of 45 calculators. The store paid $11.00 for each scientific calculator. The others all graphing calculators, cost the store $55 each. How many of each type of calculator was ordered?
Answer by mananth(16946)

total number of calculators =45
Let number of scientific calculators be x
other calculators will be 45-x
..
scientific calculator cost = $11
graphing calculator costs $ 55
..
11x+55(45-x)= 1683
11x+2475-55x=1683
-44x=1683-2475
-44x=-792
x= 18 scientific calculators
45-18= 27 all graphic calculators